OpenStudy (anonymous):

Select the correct reasons for statements 1 and 3 to complete the proof. Prove: 5x + 9 • x • 3 = 32x 1. 5x + 9 • x • 3 = 5x + 9 • 3 • x 1. ? 2. 5x + 9 • x • 3 = 5x + 27x 2. Multiplication 3. 5x + 9 • x • 3 = 32x 3. ? A. Associative Property of Addition, Addition B. Associative Property of Multiplication, Addition C. Commutative Property of Multiplication, Addition D. Commutative Property of Multiplication, Multiplication Last question

OpenStudy (anonymous):

C. Commutative Property of Multiplication, Addition

OpenStudy (anonymous):

Okay thank you ^^!

OpenStudy (anonymous):

9*x*3 became 9*3*x. a*b=b*a, 4*2=8

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

Yeah I see it now ^^

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Np